OpenStudy (aaronandyson):

A ray of light is incident at an angle of 50 degrees with the surface of glass of refractive index of 1.5.Find the deviation of ray when it enters glass.

OpenStudy (will.h):

angle of deviation in case of refraction if d=i-r you know that right?

OpenStudy (will.h):

So anyway we have i = 50 but we don't know the value or r refractive index = sin(50)/sin(r) = 1.5 sin(r) = sin(50)/1.5 = 0.5107

OpenStudy (will.h):

i got the answer 30.7 degrees check it back again am not so sure

OpenStudy (aaronandyson):

the ans is : 14 degs 38 secs

OpenStudy (aaronandyson):

idk how'd they get that

OpenStudy (alekos):

You have an angle mix up

OpenStudy (aaronandyson):

What's the mix up?

OpenStudy (alekos):

The angle of incidence is 40 degrees

OpenStudy (alekos):

That's what you need to use in the snells law equation

OpenStudy (aaronandyson):

I think the book made a typo.

OpenStudy (alekos):

No. The book is correct because I got the same answer 14.5 degrees

OpenStudy (aaronandyson):

I copied the question word for word... please show me your work..?

OpenStudy (mrnood):

 You have misread the question It says it makes an angle of 50 WITH SURFACE Therefore the angle of incidence is 40. The deviation is then 14.63° = 14 deg 38min